The Stewart Company has $2,421,000 in current assets and $1,041,030 in current liabilities. Its initial inventory level is $605,250, and it will raise funds as additional notes payable and use them to increase inventory. How much can its short-term debt (notes payable) increase without pushing its current ratio below 2.0? Round your answer to the nearest cent.
Answer #1

Current ratio = Current Assets / Current Liabilities

Let X be the Addition short tern loan that is used for Inventory.

2 = [ 2421000 + X ] / [ 1041030 + X]

2421000 + X = 2082060 +2X

2X - X = 2421000 - 2082060

X = 338940

Short term Debt can be borrowed for Inventory is $ 338,940